These problems dont make sense to me:
4(x-4)-2(x+3)≥ 2x-10

Answer by Alan3354(69443)

4(x-4)-2(x+3) >= 2x-10
Just organize and simplify.
4x-16 - (2x+6) >= 2x-10
2x - 22 >= 2x - 10
-22 >= -10
Not possible --> no solution.
ie, there is no value of x that fits. That is the solution.